# Improved Constructions and Lower Bounds for Maximally Recoverable Grid Codes

### v1

@@ I-A Literature Overview
Another well-studied regime is the case in which $h=0$. In that case, MR $(m,n,a,b,0)$ grid codes are more commonly referred to as MR $(m,n,a,b)$ tensor codes, because the generator matrix of a $(m,n,a,b)$ tensor code is the tensor (or Kronecker) product of two smaller generator matrices (corresponding to the row and column codes). This topology (although not MR) has been used in practice by Meta [2]. The case $a=1$ is the most well-studied, especially in the connection to higher order MDS codes, [17, 18, 19, 20, 21, 22, 23, 24] and in particular has been found to be closely linked to the construction of optimal list-decodable codes [20, 23, 25, 26, 27, 28]. See Brakensiek et al. [29] for an overview of field size bounds for higher order MDS codes. In general for $a,b\geq 2$ (and $h=0$) the situation is poorly understood. In fact, we even lack an efficient description of which patterns $E$ can be recovered by an MR $(m,n,2,2)$ grid code. However, Brakensiek et al. [30] recently proved that any such characterization is equivalent to characterizing the which graphs are bipartitely rigid [31]. See Brakensiek et al. [30] and references therein for a more detailed discussion.
@@ I-B Results
We also improve the lower bound (1) by showing that $q(m,n,1,1,h)\geq q(m-2,n-h,1,1,1)$, which shows that the constructing codes for $h\geq 2$ is at least as hard as for $h=1$ (see Theorem 24). We remark that some of our techniques also work in the regime in which $m$ is growing as a function of $n$. These extensions are discussed in the Appendix.
@@ II Problem Setting
To work toward these results, we now formally define the MR grid code problem. We define $m]:=\{1,2,\ldots,m\}$. In our $m\times n$ grid code, we always assume $m\leq n$. We describe a $(m,n,a=1,b=1,h)$ grid code by its parity check matrix $H\in\mathbb{F}_{q}^{(m+n+h-1)\times mn}$. Note that any $(i,j)\in[m]\times[n]$ corresponds to the column indexed n(i1)+jn(i-1)+j of $H$. In other words, column $\ell$ of $H$ corresponds to the grid square $(\lceil\ell/n\rceil,(\ell\!\!\mod n)+1)$. We now describe $H$ as follows, see Figure 1 for an example.
@@
We let $C$ denote the set of all $c\in\mathbb{F}^{mn}$ such that $Hc=0$. Given a pattern $E\subseteq[m]\times[n]$, we let $H|_{E}$ denote the restriction of $H$ to the columns indexed by $\{n(i-1)+j\mid(i,j)\in E\}$. We say that $E$ is correctable if $\operatorname{rank}H|_{E}=|E|$. In other words, for any $c\in C$ we have that $c|_{\bar{E}}=0$ iff $c=0$.
@@ II-A Correctable Patterns
We note they prove an analogous result for arbitrary $a,b\geq 1$, but we only use the $a=b=1$ setting in this paper. We say our parity check matrix $H\in\mathbb{F}_{q}^{(m+n+h-1)\times mn}$ is MR $(m,n,a=1,b=1,h)$ if $\operatorname{rank}H|_{E}=|E|$ for every pattern $E$ which is an cyclic graph plus at most $h$ edges. This criterion can be a bit cumbersome to check directly, so we next give a somewhat simpler criterion for checking if a parity check matrix is MR.
@@ II-B Cycle Sum Independence
Given, a spanning tree $T\subseteq[m]\times[n]$ (i.e., an acyclic graph for which every vertex is incident with at least one edge), for every $(i,j)\in[m]\times[n]\setminus T$, there exists a unique cycle containing $(i,j)$ and otherwise uses edges from $T$. We denote this cycle by $C_{T,(i,j)}$.
@@
By Theorem 4, we have that $H$ is MR if and only if $\operatorname{rank}H|_{E}=|E|=m+n+h-1$ for every $E$ formed by taking a spanning tree and adding $h$ additional edges. In other words, we may assume that $E=T\cup\{(i_{1},j_{1}),\ldots,(i_{h},j_{h})\}$ where $T$ is a spanning tree. Since $H$ has m+n+h1m+n+h-1 rows, it is equivalent to check that $\det H|_{E}\neq 0$.
@@
where $A_{T}$ is a $n+m-1\times n+m-1$ matrix and $B_{T}$ is a $h\times n+m-1$ matrix. Note that $\Sigma_{H}(C_{1})\cdots\Sigma_{H}(C_{h})$ form a $h\times h$ matrix, so we have that
@@
Recall we are trying to prove that $\det(H|_{E})\neq 0$ if and only if $\det(\Sigma_{H}(C_{1})\cdots\Sigma_{H}(C_{h}))$. Thus, it suffices to prove that $\det(A_{T})\neq 0$ unconditionally. To see why, note that the first m+n1m+n-1 rows of $H$ are precisely the parity check matrix of a MR $(m,n,a=1,b=1,h=0)$ tensor code. Furthermore, a spanning tree $T$ is recoverable by such an MR tensor code. Thus, $\operatorname{rank}A_{T}=|T|$. Equivalently, $\det(A_{T})\neq 0$ since $A_{T}$ is square, as desired. Therefore, $H|_{E}$ has full column rank if and only if $\Sigma_{H}(C_{T,(i_{1},j_{1})}),\ldots,\Sigma_{H}(C_{T,(i_{h},j_{h})})\in\mathbb{F}_{q}^{h}$ are linearly independent. ∎
@@
Let $E^{\prime}=C_{1}\cup\cdots\cup C_{h^{\prime}}$. By property (1), we have that $|E^{\prime}\setminus T|\leq h$ and $T$ is acyclic. Thus, we have that $E^{\prime}$ is correctable by Theorem 4. Therefore, $\operatorname{rank}H|_{E^{\prime}}=|E|$.
@@ III-A The Gabidulin Construction
The proof of Theorem 4 is constructive in the following sense. The use of Gabidulin codes from [21] converts any MR $(m,n,a,b,h=0)$ code over $\mathbb{F}_{q}$ into an MR $(m,n,a,b,h)$ code over a field of size $q^{(m-a)(n-b)}$. This implies a $2^{(m-1)(n-1)}$ field size construction for an MR $(m,n,a=1,b=1,h)$ grid code. In paper, we construct Gabidulin codes over much smaller field sizes with the same MR property.
@@
We say that our parity check matrix $H\in\mathbb{F}^{m+n+h-1\times mn}_{q}$ is a Gabidulin construction if for all $(i,j)\in[m]\times[n]$, there exists $\gamma_{i,j}\in\mathbb{F}_{q}$ such that $c^{i,j}=(\gamma_{i,j},\gamma_{i,j}^{p},\ldots,\gamma_{i,j}^{p^{h-1}}).$
Note that the map $x\mapsto x^{p}$ is a Frobenius automorphism of $\mathbb{F}_{q}$ into itself. In particular, $(x+y)^{p}=x^{p}+y^{p}$ and $(cx)^{p}=cx^{p}$ for all $x,y,c\in\mathbb{F}_{q}$. As such, for a given cycle $C=\{(i_{1},j_{1}),(i_{2},j_{1}),\ldots,(i_{k},j_{k}),(i_{1},j_{k})\}$ $\subseteq[m]\times[n]$, if $H$ is a Gabidulin construction, then $\Sigma_{H}(C)=\sum_{\ell=1}^{k}c^{i_{\ell},j_{\ell}}-c^{i_{\ell+1},j_{\ell}}=(\gamma_{C},\gamma_{C}^{p},\ldots,\gamma_{C}^{p^{h-1}}),$ where
$\gamma_{C}:=\sum_{\ell=1}^{k}\gamma_{i_{\ell},j_{\ell}}-\gamma_{i_{\ell+1},j_{\ell}}.$
@@
For all nonempty $S\subseteq[m]$, let $T_{S}\subseteq T$ be the minimal subset of $T$ connecting the left vertices indexed by $S$. We prove by induction on $S$ that $|T_{S}|\leq 2|S|-2$. The base case of $|S|=1$ is trivial, as an isolated vertex needs zero edges to be connected. Assume $|S|\geq 2$. For some $i\in S$, the vertex $i$ is incident with at most $1$ edge in $T_{S}$. Since $m]\times[n]$ is bipartite, there is a path of length two connecting $i$ to some other $j\in S\setminus\{i\}$. By the induction hypothesis, $|T_{S\setminus\{i\}}|\leq 2|S|-4$. Therefore, $|T_{S}|\leq|T_{S\setminus\{i\}}|+2\leq 2|S|-2$ as desired.
We now prove (3). For each $\ell\in[h]$, we claim that $|C_{T,(i_{\ell},j_{\ell})}\setminus T_{[m]}|\leq 2$. To see why, $j_{\ell}$ is connected to two left vertices $i_{\ell}$ and some $i^{\prime}_{\ell}\in[m]$. Since $C_{T,(i_{\ell},j_{\ell})}\setminus\{(i_{\ell},j_{\ell})\}\subseteq T$, we have that the path connecting $i_{\ell}$ and $i^{\prime}_{\ell}$ inside $C_{T,(i_{\ell},j_{\ell})}$ which avoid $j_{\ell}$ must be entirely in $T_{[m]}$. Therefore, $|C_{T,(i_{\ell},j_{\ell})}\setminus T_{[m]}|\leq|\{(i_{\ell},j_{\ell}),(i^{\prime}_{\ell},j_{\ell})\}|\leq 2$. Thus,
@@
For each $\ell\in[h]$, let $(i_{\ell},j_{\ell})=(0,m+\ell-1)$. It is straightforward to check that (3) is tight in this case.
@@ III-B Binary Encoding Construction
Let $d:=m\log_{2}n$ and let $q=2^{d}$. Note that the additive structure of $\mathbb{F}_{q}$ is isomorphic to $\mathbb{F}^{d}_{2}$. For all $(i,j)\in[m]\times[n]$, let $\gamma(i,j)\in\mathbb{F}^{d}_{2}$ have the following structure. If $i\leq m-1$, then we let $(\gamma(i,j)_{(i-1)\log_{2}n+1},\ldots,\gamma(i,j)_{i\log_{2}n})$ be the binary representation of $j$, with all other coordinates of $\gamma(i,j)$ equal to zero. Otherwise, if $i=m$, we let $\gamma(i,j)=0$.
Since we are working in characteristic two and $h=1$, it suffices to prove for every simple cycle $C\subseteq[m]\times[n]$ that $\gamma(C)\neq 0$. To see why $\gamma(C)\neq 0$, note that $C$ must have length at least 4. Thus, there are at least two left vertices $i,i^{\prime}\in[m]$ with nonzero degree in $C$. We may assume without loss of generality that $i\leq m-1$. There are exactly two distinct $j,j^{\prime}\in[n]$ for which $(i,j),(i,j^{\prime})\in C$. In particular, the first $\log_{2}n$ coordinates of $\gamma(C)$ are the XOR of the binary representations of $j$ and $j^{\prime}$. Since $j\neq j^{\prime}$, this XOR is nonzero, as desired. ∎
@@ III-D Additive Combinatorics Construction for $m=3$ and $h=1$
$\gamma(0,j)=0,\gamma(1,j)=a_{j},\gamma(2,j)=2a_{j}\text{ for all $j\in[n]$}.$
@@
because $i_{1}\neq i_{2}$ and $a_{j_{1}}\neq a_{j_{2}})$. Note this uses the fact that $q$ is odd so that $1,2,3$ are distinct in $\mathbb{F}_{q}$.

### v2

@@ I-A Literature Overview
Another well-studied regime is the case in which $h=0$. In that case, MR $(m,n,a,b,0)$ grid codes are more commonly referred to as MR $(m,n,a,b)$ tensor codes, because the generator matrix of a $(m,n,a,b)$ tensor code is the tensor (or Kronecker) product of two smaller generator matrices (corresponding to the row and column codes). This topology (although not MR) has been used in practice by Meta [2]. The case $a=1$ is the most well-studied, especially in the connection to higher order MDS codes, [17, 18, 19, 20, 21, 22, 23, 24] and in particular has been found to be closely linked to the construction of optimal list-decodable codes [20, 23, 25, 26, 27, 28, 29]. See Brakensiek et al. [30] for an overview of field size bounds for higher order MDS codes. In general for $a,b\geq 2$ (and $h=0$) the situation is poorly understood. In fact, we even lack an efficient description of which patterns $E$ can be recovered by an MR $(m,n,2,2)$ grid code. However, Brakensiek et al. [31] recently proved that any such characterization is equivalent to characterizing the which graphs are bipartitely rigid [32]. See Brakensiek et al. [31] and references therein for a more detailed discussion.
@@ I-B Results
We also improve the lower bound (1) by showing that $q(m,n,1,1,h)\geq q(m-2,n-h,1,1,1)$, which shows that the constructing codes for $h\geq 2$ is at least as hard as for $h=1$ (see Theorem 26). We remark that some of our techniques also work in the regime in which $m$ is growing as a function of $n$. These extensions are discussed in the Appendix, including the following result:
Theorem 4 (Informal version of Theorem 30).
Assume that $m\leq n$ are powers of two, then $q(m,n,1,1,1)\leq 8(8n/m)^{m-1}$.
@@ II Problem Setting
To work toward these results, we now formally define the MR grid code problem. We define $m]:=\{1,2,\ldots,m\}$. In our $m\times n$ grid code, we always assume $m\leq n$. We describe a $(m,n,a=1,b=1,h)$ grid code by its parity check matrix $H\in\mathbb{F}_{q}^{(m+n+h-1)\times mn}$. For this parity check matrix to be sensible, we always assume that $m+n+h-1\leq mn$. Note that any $(i,j)\in[m]\times[n]$ corresponds to the column indexed n(i1)+jn(i-1)+j of $H$. In other words, column $\ell$ of $H$ corresponds to the grid square $(\lceil\ell/n\rceil,(\ell\!\!\mod n)+1)$. We now describe $H$ as follows, see Figure 1 for an example.
@@
We let $C$ denote the set of all $c\in\mathbb{F}^{mn}$ such that $Hc=0$. Given a pattern $E\subseteq[m]\times[n]$, we let $H|_{E}$ denote the restriction of $H$ to the columns indexed by $\{n(i-1)+j\mid(i,j)\in E\}$. We say that $E$ is correctable for any $c\in C$ we have that $c|_{\bar{E}}=0$ iff $c=0$, where $\bar{E}=[m]\times[n]\setminus E$. Equivalently, $E$ is not correctable if there exists a nonzero $c\in C$ such that $c_{i,j}\neq 0$ only if $(i,j)\in E$. Since $Hc=0$ for all $c\in C$, the existence of such a codeword is equivalent to there being a linear dependence among the columns of $H|_{E}$. In other words, we have the following proposition.
Proposition 5.
$E\subseteq[m]\times[n]$ is correctable if and only if $\operatorname{rank}H|_{E}=|E|$.
@@ II-A Correctable Patterns
We note they prove an analogous result for arbitrary $a,b\geq 1$, but we only use the $a=b=1$ setting in this paper. We say our parity check matrix $H\in\mathbb{F}_{q}^{(m+n+h-1)\times mn}$ is MR $(m,n,a=1,b=1,h)$ if $\operatorname{rank}H|_{E}=|E|$ for every pattern $E$ which is an acyclic graph plus at most $h$ edges. This criterion can be a bit cumbersome to check directly, so we next give a somewhat simpler criterion for checking if a parity check matrix is MR.
@@ II-B Cycle Sum Independence
Consider a spanning tree $T\subseteq[m]\times[n]$ (i.e., an acyclic graph for which every vertex is incident with at least one edge). For every $(i,j)\in[m]\times[n]\setminus T$ note that there is a unique path connecting $i$ and $j$. Since $i$ and $j$ are on opposite sides of a bipartite graph, this path has an odd number of edges, say 2k12k-1. Denote the vertices of this path by $i=p_{1},p_{2},\ldots,p_{2k}=j$. Since this path from $i$ to $j$ within $T$ is unique, there is exactly one cycle which uses edges from $T\cup\{(i,j)\}$. Precisely, this cycle is
$C_{T,(i,j)}:=\{(i,j),(i=p_{1},p_{2}),(p_{3},p_{2}),(p_{3},p_{4})\ldots,(p_{2k-1},p_{2k-2}),(p_{2k-1},j=p_{2k})\}.$
For example, if $m=n=2$, and $T=\{(1,1),(1,2),(2,2)\}$, then $(2,1)\in[m]\times[n]\setminus T$ and $C_{T,(2,1)}=\{(2,1),(2,2),(1,2),(1,1)\}$.
@@
By Theorem 6, we have that $H$ is MR if and only if $\operatorname{rank}H|_{E}=|E|$ for every $E$ which is the union of an acyclic graph and at most $h$ additional edges. If there is a strictly larger $E^{\prime}\supsetneq E$ of this form, then note that $\operatorname{rank}H|_{E^{\prime}}=|E^{\prime}|$ implies $\operatorname{rank}H|_{E}=|E|.$ In other words, we may assume that $E=T\cup\{(i_{1},j_{1}),\ldots,(i_{h},j_{h})\}$ where $T$ is a spanning tree [^1] If $E$ does not contain a spanning tree, then viewed as a subgraph of $m]\times[n]$, $E$ has multiple connected components. Adding any edge bridging these connected components will not change the property that $E$ is an acyclic graph plus at most $h$ additional edges. and $(i_{1},j_{1}),\ldots,(i_{h},j_{h})\not\in T$ are distinct. Therefore, |E|=m+n1+h|E|=m+n-1+h. Since $H$ has m+n+h1m+n+h-1 rows, it is equivalent to check that $\det H|_{E}\neq 0$.
@@
where $A_{T}$ is a $(n+m-1)\times(n+m-1)$ matrix and $B_{T}$ is a $h\times(n+m-1)$ matrix. Note that $\Sigma_{H}(C_{1})\cdots\Sigma_{H}(C_{h})$ form a $h\times h$ matrix, so we have that
@@
Recall we are trying to prove that $\det(H|_{E})\neq 0$ if and only if $\det(\Sigma_{H}(C_{1})\cdots\Sigma_{H}(C_{h}))$. Thus, it suffices to prove that $\det(A_{T})\neq 0$ unconditionally. To see why, note that the first m+n1m+n-1 rows of $H$ are precisely the parity check matrix of a MR $(m,n,a=1,b=1,h=0)$ tensor code. Furthermore, the erasure pattern corresponding to the edges of a spanning tree $T$ is recoverable by such an MR tensor code [1]. Thus, $\operatorname{rank}A_{T}=|T|$. Equivalently, $\det(A_{T})\neq 0$ since $A_{T}$ is square, as desired. Therefore, $H|_{E}$ has full column rank if and only if $\Sigma_{H}(C_{T,(i_{1},j_{1})}),\ldots,\Sigma_{H}(C_{T,(i_{h},j_{h})})\in\mathbb{F}_{q}^{h}$ are linearly independent. ∎
@@
Let $E^{\prime}=C_{1}\cup\cdots\cup C_{h^{\prime}}$. By property (1), we have that $|E^{\prime}\setminus T|\leq h$ and $T$ is acyclic. Thus, we have that $E^{\prime}$ is correctable by Theorem 6. Therefore, $\operatorname{rank}H|_{E^{\prime}}=|E^{\prime}|$.
@@ III-A The Gabidulin Construction
The proof of Theorem 6 is constructive in the following sense. The use of Gabidulin codes from [21] converts any MR $(m,n,a,b,h=0)$ code over $\mathbb{F}_{q}$ into an MR $(m,n,a,b,h)$ code over a field of size $q^{(m-a)(n-b)}$. Since an MR $(m,n,a=1,b=1,h=0)$ code exists over $\mathbb{F}_{2}$ (see [1]), we obtain a $2^{(m-1)(n-1)}$ field size construction for an MR $(m,n,a=1,b=1,h)$ grid code. In paper, we construct Gabidulin codes over much smaller field sizes with the same MR property.
@@
We say that our parity check matrix $H\in\mathbb{F}^{m+n+h-1\times mn}_{q}$ is a Gabidulin construction if for all $(i,j)\in[m]\times[n]$, there exists $\gamma(i,j)\in\mathbb{F}_{q}$ such that $c^{i,j}=(\gamma(i,j),\gamma(i,j)^{p},\ldots,\gamma(i,j)^{p^{h-1}}).$
Note that the map $x\mapsto x^{p}$ is a Frobenius automorphism of $\mathbb{F}_{q}$ into itself. In particular, $(x+y)^{p}=x^{p}+y^{p}$ and $(cx)^{p}=cx^{p}$ for all $x,y\in\mathbb{F}_{q}$ and $c\in\mathbb{F}_{p}$, where $\mathbb{F}_{p}$ is the base subfield of $\mathbb{F}_{q}$. As such, for a given cycle $C=\{(i_{1},j_{1}),(i_{2},j_{1}),\ldots,(i_{k},j_{k}),(i_{1},j_{k})\}$ $\subseteq[m]\times[n]$, if $H$ is a Gabidulin construction, then $\Sigma_{H}(C)=\sum_{\ell=1}^{k}c^{i_{\ell},j_{\ell}}-c^{i_{\ell+1},j_{\ell}}=$ $(\gamma(C),\gamma(C)^{p},$ $\ldots,\gamma(C)^{p^{h-1}}),$ where
$\gamma(C):=\sum_{\ell=1}^{k}\gamma(i_{\ell},j_{\ell})-\gamma(i_{\ell+1},j_{\ell}).$
@@
If $m=1$, then $T=[m]\times[n]$ so the LHS is (3) is always zero, while $2(m+h-1)=2h\geq 0$. Thus, we assume $m\geq 2$.
Say an edge $(i,j)\in T$ is right-lonely if $(i,j)$ is the only edge of $T$ containing $j\in[n]$. Let $T^{\prime}\subseteq T$ be the (maximal) subgraph of $T$ with all right-lonely edges removed. Let $S\subseteq[n]$ be the right vertices incident to $T^{\prime}$. As such, $T^{\prime}$ has at most |S|+m1|S|+m-1 edges. Thus, in order for $T^{\prime}$ to lack any further right-lonely edges, we must have that $|S|+m-1\geq|T^{\prime}|\geq 2|S|$. Therefore, $m-1\geq|S|$, so $|T^{\prime}|\leq 2(m-1)$.
We now prove (3). For each $\ell\in[h]$, we claim that $|C_{T,(i_{\ell},j_{\ell})}\setminus T^{\prime}|\leq 2$. To see why, $j_{\ell}$ is connected to two left vertices $i_{\ell}$ and some $i^{\prime}_{\ell}\in[m]$. Since $C_{T,(i_{\ell},j_{\ell})}\setminus\{(i_{\ell},j_{\ell})\}\subseteq T$, we have that the path connecting $i_{\ell}$ and $i^{\prime}_{\ell}$ inside $C_{T,(i_{\ell},j_{\ell})}$ which avoids $j_{\ell}$ must have that every right vertex has degree at least two. Therefore, $C_{T,(i_{\ell},j_{\ell})}\setminus\{(i_{\ell},j_{\ell}),(i^{\prime}_{\ell},j_{\ell})\subseteq T^{\prime}$. Hence, $|C_{T,(i_{\ell},j_{\ell})}\setminus T^{\prime}|\leq|\{(i_{\ell},j_{\ell}),(i^{\prime}_{\ell},j_{\ell})\}|\leq 2$. Thus,
@@
For each $\ell\in[h]$, let $(i_{\ell},j_{\ell})=(1,m+\ell-1)$. It is straightforward to check that (3) is tight in this case.
@@ III-B Binary Encoding Construction
Let $d:=(m-1)\log_{2}n$ and let $q=2^{d}$. Note that the additive structure of $\mathbb{F}_{q}$ is isomorphic to $\mathbb{F}^{d}_{2}$. For all $(i,j)\in[m]\times[n]$, let $\gamma(i,j)\in\mathbb{F}^{d}_{2}$ have the following structure. If $i\leq m-1$, then we let $(\gamma(i,j)_{(i-1)\log_{2}n+1},\ldots,\gamma(i,j)_{i\log_{2}n})$ be the binary representation of $j$, with all other coordinates of $\gamma(i,j)$ equal to zero. Otherwise, if $i=m$, we let $\gamma(i,j)=0$.
Since we are working in characteristic two and $h=1$, it suffices to prove for every simple cycle $C\subseteq[m]\times[n]$ that $\gamma(C)\neq 0$. To see why $\gamma(C)\neq 0$, note that $C$ must have length at least 4. Thus, there are at least two left vertices $i,i^{\prime}\in[m]$ with nonzero degree in $C$. We may assume without loss of generality that $i\leq m-1$. There are exactly two distinct $j,j^{\prime}\in[n]$ for which $(i,j),(i,j^{\prime})\in C$. In particular, the coordinates of $\gamma(C)$ indexed by $\{(i-1)\log_{2}n+1,\ldots,i\log_{2}n\}$ are the XOR of the binary representations of $j$ and $j^{\prime}$. Since $j\neq j^{\prime}$, this XOR is nonzero, as desired. ∎
@@ III-D Additive Combinatorics Construction for $m=3$ and $h=1$
$\gamma(1,j)=0,\gamma(2,j)=a_{j},\gamma(3,j)=2a_{j}\text{ for all $j\in[n]$}.$
@@
because $i_{1}\neq i_{2}$ and $a_{j_{1}}\neq a_{j_{2}}$. Note this uses the fact that $q$ is odd so that $1,2,3$ are distinct in $\mathbb{F}_{q}$.